OpenStudy (anonymous):

Ill give a medal if you can solve this Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = −2 x = 0 x = 2 There is no x value, as there is no solution. x can be any value, as there is an infinite number of solutions.

OpenStudy (anonymous):

You have a System of linear equations. So to solve this, using substitution, then you need to isolate x or y in the first equation. And insert the value into the 2nd equation. \[2x+y=1<=>y=1−2x\] Now I insert the value for y: \[4x+2(1−2x)=−2\] Now you have one equation with one variable. Can you solve x from here?